# Exploring the Potential of Large Language Models for Heterophilic Graphs

### v2

@@ Abstract
Yuxia Wu1, Shujie Li2, Yuan Fang1, Chuan Shi2 1Singapore Management University 2Beijing University of Post and Telecommunication yieshah2017@gmail.com, shujieli@bupt.edu.cn yfang@smu.edu.sg, shichuan@bupt.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of applications, from natural language processing Brown et al. (2020) to computer vision Wang et al. (2024b), leveraging the extensive open-world knowledge LLMs encode. Inspired by these successes, recent efforts have extended the application of LLMs to the graph domain, particularly in text-attributed graphs where node attributes are composed of textual information Li et al. (2023); Zhang et al. (2024). In the context of heterophilic graphs, where connected nodes often exhibit contrasting features or class labels Bo et al. (2021); Song et al. (2023); Liang et al. (2024), LLMs offer a unique opportunity to enhance the understanding of complex semantic relationships between these connected nodes, which remains largely unexplored.
Existing approaches to addressing heterophily in GNNs typically involve extracting shallow embeddings from textual information, using them as initial node features without fully exploiting their rich semantic content. They can be broadly categorized into two main strategies: non-local neighbor extension and architectural refinement Zheng et al. (2022); Gong et al. (2024). The former extends the node’s receptive field to include distant, high-order neighbors Abu-El-Haija et al. (2019); Song et al. (2023) or potential connections Jin et al. (2021); Wang and Zhang (2022); Zou et al. (2023), thereby enhancing node representations through a broader scope of information integration. The latter modifies the core functions of GNNs, such as the message aggregation and updating functions, to better suit heterophilic contexts.
In summary, current methods for heterophilic graphs largely overlook the rich textual content associated with the nodes in real-world graphs, which can provide deeper insights into heterophilic contexts. For instance, in a website, the textual content on hyperlinked webpages can enrich the understanding and prediction of heterophilic links. While traditional text processing techniques used in GNNs, such as bag-of-words or shallow embeddings, are inadequate for capturing complex semantics, LLMs Zhao et al. (2023b) have extended their applications to empower GNNs for text-attributed graphs Liu et al. (2023); Li et al. (2023); Yu et al. (2024); Mao et al. (2024). However, these efforts focus on homophilic graphs, leaving heterophilic graphs largely unexplored.
@@
Additionally, LLMs often incur high computational costs even for inference, limiting their practical deployment for real-world applications. To this end, we further explore model distillation techniques Xu et al. (2024) to condense the knowledge from fine-tuned LLMs into small language models (SLMs) Schick and Schütze (2021); Gu et al. (2024) to speed up the inference stage required for edge discrimination and reweighting. Specifically, we utilize the LLM well-tuned for edge discrimination to generate high-quality pseudo-labels for heterophilic and homophilic edges. These pseudo-labels supplement the limited ground truth labels, forming an expanded label set that further enables the fine-tuning of SLMs. The fine-tuned SLM then replaces the LLM for conducting inference for the edge discrimination and reweighting process, maintaining effectiveness while significantly reducing inference time.
@@ 2 Related Work
LLMs for graph learning.
The existing research about LLMs for graph learning includes LLM-based methods that adopt LLM as the backbone and GNN+LLM-based methods that integrate the advantages of both GNNs and LLMs Liu et al. (2023); Chen et al. (2024). The former works focus on aligning graph data with natural language via graph-to-token and graph-to-text approaches Liu et al. (2023). The graph-to-token approach involves tokenizing graph data to align it with natural language, enabling joint understanding with data from other modalities Zhao et al. (2023a); Ye et al. (2024). Graph-to-text focuses on describing graph information using natural language Liu and Wu (2023); Wang et al. (2024a); Guo et al. (2023). The latter harnesses the strengths of both language understanding from LLMs and structural analysis from GNNs by using GNN-centric methods utilizing LLMs to extract node features from raw data and make predictions using GNNs He et al. (2024); Xie et al. (2023) or LLM-centric methods utilizing GNNs to enhance the performance of LLM Tang et al. (2024); Zhang et al. (2024).
Heterophilic graph learning.
The existing approaches generally fall into two main strategies: non-local neighbor extension and architectural refinement Zheng et al. (2022); Gong et al. (2024). Non-local neighbor extension approaches aim to extend the neighbors to include non-local nodes in the graph that may share similar labels or features. These methods often involve high-order neighbor mixing Abu-El-Haija et al. (2019); Song et al. (2023) or discovering potential neighbors based on various distance measurements, such as feature-based distance Jin et al. (2021); Bodnar et al. (2022), structure-based distance Pei et al. (2020), or hybrid approaches Zou et al. (2023); Wang and Zhang (2022); Li et al. (2022). Architectural refinement approaches enhance the GNN architecture by employing identifiable message aggregation to discriminate and amplify messages from similar neighbors while minimizing the influence of dissimilar ones Bo et al. (2021); Zhu et al. (2021); Liang et al. (2024), or by leveraging inter-layer combinations to capture information from different neighbor ranges Xu et al. (2018); Chien et al. (2021); Zhu et al. (2020), thereby improving the model’s representation power in heterophilic graphs.
@@ 3.1 Preliminaries
Problem formulation.
Let $G=(V,E,X,C)$ denotes a text-attributed graph with a set of nodes $V$ and a set of edges $E$, where each node $v\in V$ is associated with a text document $x_{v}\in X$. $C$ is the set of node classes. In this paper, we address the task of semi-supervised transductive node classification for heterophilic graphs. Specifically, a subset of the nodes is designated as the training/validation nodes with known class labels, while the goal is to predict the unknown labels of the remaining nodes in the graph.
Classic GNNs.
GNNs typically employ a multi-layer approach to neighborhood aggregation, wherein each node incrementally gathers and aggregates contexts from its neighboring nodes. At the $l^{\text{th}}$ layer, the representation $\mathbf{h}_{v}^{l}\in\mathbb{R}^{d_{l}}$ of a node $v$ is derived as follows:
$\mathbf{h}_{v}^{(l)}=\sigma(\text{AGGR}(\mathbf{h}_{v}^{(l-1)},\{\mathbf{h}_{u}^{(l-1)}\mid u\in\mathcal{N}(v)\})),$ (1)
where $d_{l}$ is the dimension of the node representations at the $l^{\text{th}}$ layer. The function $\text{AGGR}(\cdot)$ denotes an aggregation mechanism combining the feature vectors of the neighboring nodes, where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. $\sigma$ represents an activation function.
@@ 3.2 Overall Framework
Figure 1 illustrates the overall framework of our approach LLM4HeG, with a two-stage framework leveraging LLMs for heterophilic graph modeling. As shown in Fig. 1(b), Stage 1 involves our LLM-enhanced edge discriminator, where we fine-tune an LLM to distinguish between homophilic and heterophilic edges, utilizing the rich textual data associated with nodes and a limited amount of ground truth label. Following this, in Fig. 1(c), Stage 2 involves LLM-guided edge reweighting to learn adaptive weights for both homophilic and heterophilic edges. These weights are adapted to individual edges by integrating node features, graph structures, and edge types, enabling fine-grained aggregation within GNNs.
@@
3.3 Stage 1: LLM-enhanced Edge Discriminator
@@
Fine-tuning.
The template, together with the ground-truth labels, enables the fine-tuning of LLMs. For efficiency, we adopt a parameter-efficient fine-tuning technique called LoRA, which strategically updates only a small fraction of the LLM’s parameters Hu et al. (2022).
Inference.
After fine-tuning the LLM for edge discrimination, we can employ it to infer the relationship between any two nodes on the graph, determining whether they have a potentially heterophilic or homophilic edge. During inference, we use the same template to generate input text for each node pair, which is then fed into the fine-tuned LLM to produce a “Yes” or “No” answer regarding the edge type.
3.4 Stage 2: LLM-guided Edge Reweighting
@@
$w_{uv}^{\text{LLM}}$ =$=\begin{cases}\tanh(w_{\text{Ho}})&\text{if }O_{\text{LLM}}(u,v)=\textit{Yes},\\ \tanh(w_{\text{He}})&\text{if }O_{\text{LLM}}(u,v)=\textit{No}\end{cases}$ (3)
@@
$w_{uv}^{\text{G}}=\tanh\left(\mathbf{g}^{\top}\left[\mathbf{h}_{u}\parallel\mathbf{h}_{v}\right]\right)$ (4)
where $\parallel$ denotes the concatenation operation, and $\mathbf{g}$ is learnable, which can be seen as a shared convolutional kernel Veličković et al. (2018).
@@
$w_{uv}$ $=\frac{1}{2}\left(w_{uv}^{\text{LLM}}+w_{uv}^{\text{G}}\right).$ (5)
@@
$\mathbf{h}_{v}^{(l)}$ $=\epsilon\mathbf{h}_{v}^{(0)}+\sum_{u\in\mathcal{N}_{i}(v)}\frac{w_{uv}}{\sqrt{d_{u}d_{v}}}\mathbf{h}_{u}^{(l-1)},$ (6)
@@
Training.
We adopt a typical cross-entropy (CE) loss to train node classification. Additionally, to ensure the learned weights $w_{\text{Ho}}$ and $w_{\text{He}}$ show sufficient separation, we introduce a margin-based regularization term into the CE loss.
@@
Inference.
During the inference phase of Stage 2, we apply the LLM-enhanced edge discriminator fine-tuned in Stage 1 to generate the edge type for each node pair involved in the test set. Subsequently, we can calculate the output representations for the test nodes based on the reweighted edges for node classification.
@@ 4.1 Experimental Setup
Datasets.
Given that the datasets commonly employed in heterophily graph tasks lack original textual information, we collect publicly available raw text directly from the original data providers and preprocess these datasets. Consequently, our experiments only include datasets that contain raw text. Specifically, we have prepared five datasets: Cornell, Texas, and Wisconsin Pei et al. (2020), Actor Tang et al. (2009) and Amazon Platonov et al. (2023). The Cornell, Texas, and Wisconsin datasets Pei et al. (2020) are derived from computer science department websites where webpages serve as nodes and hyperlinks as edges. The Actor dataset Tang et al. (2009) is an actor co-occurrence network with nodes representing actors and edges denoting their collaborations. The Amazon dataset Platonov et al. (2023) is a product co-purchasing network, where nodes represent products and edges link products frequently bought together. The statistical details of the datasets are presented in Table 1 with more details provided in Appendix A. Additionally, we include the edge homophily score $\mathcal{H}(G)$ Yan et al. (2022) for each dataset, which quantifies the level of homophily (1 means perfectly homophily while 0 stands for total heterophily).
@@
Baselines.
To comprehensively evaluate the performance of the proposed model, we compare it against a set of baseline models that fall into two main categories: classic GNN models and heterophily-specific models. The classic GNN models include GCN Kipf and Welling (2016), GraphSAGE Hamilton et al. (2017) and GAT Veličković et al. (2018). The Heterophily-specific models include H2GCN Zhu et al. (2020), FAGCN Bo et al. (2021), JacobiConv Wang and Zhang (2022), OGNN Song et al. (2023), SEGSL Zou et al. (2023), DisamGCL Zhao et al. (2024) with more details provided in Appendix D.
Implementation Details.
We adopt Vicuna-v1.5-7B Zheng et al. (2024) as the LLM model and the Bloom model Le Scao et al. (2023) with 560M and 1B parameters as SLMs. Consistent with most existing works Zhu et al. (2020), we randomly split the nodes into train, validation and test sets with a proportion of 48%/32%/20% for Cornell, Texas, Wisconsin and Actor datasets. We set the proportion of Amazon dataset as 50%/25%/25% following Platonov et al. (2023). In our main experiments, we use FAGCN Bo et al. (2021) as the GNN backbone, adhering to the parameter settings outlined in FAGCN. We only fine-tune the other hyperparameters based on the performance observed on the validation sets. All experiments are repeated 10 times, and we report the averaged results with standard deviation. To ensure a fair comparison, we use the initial node features derived from the Vicuna 7B model for all methods. More implementation details and hyper-parameters are introduced in Appendix B and C.

### v3

@@ Abstract
Yuxia Wu1, Shujie Li2, Yuan Fang1, Chuan Shi2 1Singapore Management University, 2Beijing University of Post and Telecommunication yieshah2017@gmail.com, shujieli@bupt.edu.cn, yfang@smu.edu.sg, shichuan@bupt.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of applications, from natural language processing Brown et al. (2020) to computer vision Wang et al. (2024b), leveraging the extensive open-world knowledge LLMs encode. Inspired by these successes, recent efforts have extended the application of LLMs to the graph domain, particularly in text-attributed graphs where node attributes are composed of textual information Li et al. (2023b); Zhang et al. (2024). In the context of heterophilic graphs, where connected nodes often exhibit contrasting features or class labels Bo et al. (2021); Sun et al. (2022); Liang et al. (2024), LLMs offer a unique opportunity to enhance the understanding of complex semantic relationships between these connected nodes, which remains largely unexplored.
Existing approaches to addressing heterophily in GNNs typically involve extracting shallow embeddings from textual information, using them as initial node features without fully exploiting their rich semantic content. They can be broadly categorized into two main strategies: non-local neighbor extension and architectural refinement Zheng et al. (2022); Gong et al. (2024). The former extends the node’s receptive field to include distant, high-order neighbors Abu-El-Haija et al. (2019); Song et al. (2023) or potential connections Jin et al. (2021); Wang and Zhang (2022); Zou et al. (2023), thereby enhancing node representations through a broader scope of information integration. The latter modifies the core functions of GNNs, such as the message aggregation and updating functions, to better suit heterophilic contexts Du et al. (2022).
In summary, current methods for heterophilic graphs largely overlook the rich textual content associated with the nodes in real-world graphs, which can provide deeper insights into heterophilic contexts. For instance, textual content on hyperlinked webpages can enrich the understanding and prediction of heterophilic links. Traditionally, GNNs employ bag-of-words or shallow embeddings to incorporate textual attributes, which are inadequate for capturing complex semantics. While LLMs Zhao et al. (2023b) have been used to empower GNNs for text-attributed graphs Liu et al. (2023); Li et al. (2023b); Yu et al. (2024a); Mao et al. (2024), existing efforts focus on homophilic graphs, leaving heterophilic graphs largely unexplored.
@@
Additionally, LLMs often incur high computational costs even for inference, limiting their practical deployment for real-world applications. To this end, we further explore model distillation techniques Xu et al. (2024) to condense the knowledge from fine-tuned LLMs into small language models (SLMs) Schick and Schütze (2021); Li et al. (2023a); Gu et al. (2024); Pan et al. (2024) to speed up the inference stage required for edge discrimination and reweighting. Specifically, we utilize the LLM well-tuned for edge discrimination to generate high-quality pseudo-labels for heterophilic and homophilic edges. These pseudo-labels supplement the limited ground truth labels, forming an expanded label set that further enables the fine-tuning of SLMs. The fine-tuned SLM then replaces the LLM for conducting inference for the edge discrimination and reweighting, maintaining effectiveness while significantly reducing inference time.
@@ 2 Related Work
We review the literature on LLM-based and heterophilic graph learning and highlight the key distinctions of our work from existing studies.
Existing research on LLMs for graph learning includes LLM-based methods that adopt LLM as the backbone and GNN+LLM-based methods that integrate the advantages of both GNNs and LLMs Liu et al. (2023); Chen et al. (2024). The former works focus on aligning graph data with natural language via graph-to-token and graph-to-text approaches Liu et al. (2023). The graph-to-token approach involves tokenizing graph data to align it with natural language, enabling joint understanding with data from other modalities Zhao et al. (2023a); Ye et al. (2024). Graph-to-text focuses on describing graph information using natural language Liu and Wu (2023); Wang et al. (2024a); Guo et al. (2023). The latter harnesses the strengths of both language understanding from LLMs and structural analysis from GNNs by using GNN-centric methods utilizing LLMs to extract node features from raw data and make predictions using GNNs He et al. (2024); Xie et al. (2023) or LLM-centric methods utilizing GNNs to enhance the performance of LLM Tang et al. (2024); Zhang et al. (2024).
Existing heterophilic graph learning approaches generally fall into two main strategies: non-local neighbor extension and architectural refinement Zheng et al. (2022); Gong et al. (2024). Non-local neighbor extension approaches aim to extend the neighbors to include non-local nodes in the graph that may share similar labels or features. These methods often involve high-order neighbor mixing Abu-El-Haija et al. (2019); Song et al. (2023); Yu et al. (2024b) or discovering potential neighbors based on various distance measurements, such as feature-based distance Jin et al. (2021); Bodnar et al. (2022), structure-based distance Pei et al. (2020), or hybrid approaches Wang and Zhang (2022); Wang et al. (2022); Li et al. (2022); Zou et al. (2023); Bi et al. (2024). Architectural refinement approaches enhance the GNN architecture by employing identifiable message aggregation to discriminate and amplify messages from similar neighbors while minimizing the influence of dissimilar ones Bo et al. (2021); Zhu et al. (2021); Du et al. (2022); Liang et al. (2024), or by leveraging inter-layer combinations to capture information from different neighbor ranges Xu et al. (2018); Chien et al. (2021); Zhu et al. (2020), thereby improving the model’s representation power in heterophilic graphs.
The key distinctions of our method lie in two aspects. First, we explore LLMs to enhance text-attributed heterophilic graphs’ modeling specifically. While prior works leverage LLMs for text-attributed graphs, including edge reweighting Sun et al. (2023); Ling et al. (2024), they do not explicitly target heterophilic graphs. Additionally, existing heterophilic graph methods often overlook rich textual node attributes, which provide essential semantic contexts. Second, our two-stage framework employs an LLM-enhanced edge discriminator to predict edge types, followed by adaptive message propagation in GNNs using a comprehensive suite of information, including node semantics, structural contexts, and LLM-inferred edge characteristics. While GBK-GNN Du et al. (2022) follows a similar two-stage approach, it does not leverage the power of LLMs.
@@ 3.1 Preliminaries
Problem formulation. Let $G=(V,E,X,C)$ denotes a text-attributed graph with a set of nodes $V$ and a set of edges $E$, where each node $v\in V$ is associated with a text document $x_{v}\in X$. $C$ is the set of node classes. In this paper, we address the task of semi-supervised transductive node classification for heterophilic graphs. Specifically, a subset of the nodes is designated as the training/validation nodes with known class labels, while the goal is to predict the unknown labels of the remaining nodes in the graph.
Classic GNNs. GNNs typically employ a multi-layer approach to neighborhood aggregation, wherein each node incrementally gathers and aggregates contexts from its neighboring nodes. At the $l^{\text{th}}$ layer, the representation $\mathbf{h}_{v}^{l}\in\mathbb{R}^{d_{l}}$ of a node $v$ is derived as follows:
$\mathbf{h}_{v}^{(l)}=\text{AGGR}(\mathbf{h}_{v}^{(l-1)},\{\mathbf{h}_{u}^{(l-1)}:u\in\mathcal{N}(v)\}),$ (1)
where $d_{l}$ is the dimension of the node representations at the $l^{\text{th}}$ layer. The function $\text{AGGR}(\cdot)$ denotes an aggregation mechanism combining the feature vectors of the neighboring nodes, where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$.
@@ 3.2 Overall Framework
Fig. 1 illustrates the overall framework of our approach LLM4HeG, with a two-stage framework leveraging LLMs for heterophilic graph modeling. As shown in Fig. 1(b), Stage 1 involves our LLM-enhanced edge discriminator, where we fine-tune an LLM to distinguish between homophilic and heterophilic edges, utilizing the rich textual data associated with nodes and a limited amount of ground truth label. Following this, in Fig. 1(c), Stage 2 involves LLM-guided edge reweighting to learn adaptive weights for both homophilic and heterophilic edges. These weights are adapted to individual edges by integrating node features, graph structures, and edge types, enabling fine-grained aggregation within GNNs.
@@
3.3 LLM-enhanced Edge Discriminator
@@
Fine-tuning. The template, together with the ground-truth labels, enables the fine-tuning of LLMs. For efficiency, we adopt a parameter-efficient fine-tuning technique called LoRA, which strategically updates only a small fraction of the LLM’s parameters Hu et al. (2022). For fine-tuning, we use a typical cross-entropy loss to align the model output with the ground-truth responses:
$\mathcal{L}_{\text{fine-tune}}=\textstyle-\frac{1}{N}\sum_{i=1}^{N}\log P_{\theta}(x_{i}|x_{<i}),$ (2)
where $N$ is the number of tokens in the sentences, $x_{i}$ is the $i$-th token to be predicted, and $x_{<i}$ indicates previously generated tokens. $P_{\theta}(x_{i}|x_{<i})$ represents the probability of the token $x_{i}$ given the previous tokens generated by the model.
Inference. After fine-tuning the LLM for edge discrimination, we can employ it to infer the relationship between any two nodes on the graph, determining whether they have a potentially heterophilic or homophilic edge. During inference, we use the same template to generate input text for each node pair, which is then fed into the fine-tuned LLM to produce a “Yes” or “No” answer regarding the edge type.
3.4 LLM-guided Edge Reweighting
@@
$w_{uv}^{\text{LLM}}$ =$=\begin{cases}\tanh(w_{\text{Ho}})&\text{if }O_{\text{LLM}}(u,v)=\textit{Yes},\\ \tanh(w_{\text{He}})&\text{if }O_{\text{LLM}}(u,v)=\textit{No},\end{cases}$ (4)
@@
$w_{uv}^{\text{G}}=\tanh\left(\mathbf{g}^{\top}\left[\mathbf{h}_{u}\parallel\mathbf{h}_{v}\right]\right),$ (5)
where $\parallel$ denotes the concatenation operation, and $\mathbf{g}$ is a linear layer to map the concatenated feature into a scalar value, which can be seen as a shared convolutional kernel Veličković et al. (2018).
@@
$w_{uv}$ $=\textstyle\frac{1}{2}\left(w_{uv}^{\text{LLM}}+w_{uv}^{\text{G}}\right).$ (6)
@@
$\mathbf{h}_{v}^{(l)}$ $=\textstyle\epsilon\mathbf{h}_{v}^{(0)}+\sum_{u\in\mathcal{N}_{i}(v)}\frac{w_{uv}}{\sqrt{d_{u}d_{v}}}\mathbf{h}_{u}^{(l-1)},$ (7)
@@
Training. We adopt a typical cross-entropy (CE) loss to train node classification. Additionally, to ensure the learned weights $w_{\text{Ho}}$ and $w_{\text{He}}$ show sufficient separation, we introduce a margin-based regularization term into the CE loss.
@@
Inference. During the inference phase of Stage 2, we apply the LLM-enhanced edge discriminator fine-tuned in Stage 1 to generate the edge type for each node pair involved in the test set. Subsequently, we can calculate the output representations for the test nodes based on the reweighted edges for node classification.
@@ 3.5 LLMs-to-SLMs Distillation
Learning objectives. We outline the objectives of knowledge distillation in the two stages. In Stage 1, the fine-tuning of SLMs follows the same approach as fine-tuning the LLM shown in Eq. (2). In Stage 2, we train the edge reweighting using the same cross-entropy objective shown in Eq. (9).
@@ 4.1 Experimental Setup
Datasets. Given that the datasets commonly employed in heterophily graph tasks lack original textual information, we collect publicly available raw text directly from the original data providers and preprocess these datasets. Consequently, our experiments only include datasets that contain raw text. Specifically, we have prepared five datasets: Cornell, Texas, and Wisconsin Pei et al. (2020), Actor Tang et al. (2009) and Amazon Platonov et al. (2023). The Cornell, Texas, and Wisconsin datasets Pei et al. (2020) are derived from computer science department websites where webpages serve as nodes and hyperlinks as edges. The Actor dataset Tang et al. (2009) is an actor co-occurrence network with nodes representing actors and edges denoting their collaborations. The Amazon dataset Platonov et al. (2023) is a product co-purchasing network, where nodes represent products and edges link products frequently bought together. The statistical details of the datasets are presented in Table 1 with more details provided in Appendix A. Additionally, we include the edge homophily score $\mathcal{H}(G)$ Yan et al. (2022) for each dataset, which quantifies the level of homophily (1 means perfectly homophily while 0 stands for total heterophily).
@@
Baselines. We compare our method against a set of baseline models that fall into two main categories: classic GNN models and heterophily-specific models. The classic GNN models include GCN Kipf and Welling (2016), GraphSAGE Hamilton et al. (2017) and GAT Veličković et al. (2018). The heterophily-specific models include H2GCN Zhu et al. (2020), FAGCN Bo et al. (2021), JacobiConv Wang and Zhang (2022), GBK-GNN Du et al. (2022), OGNN Song et al. (2023), SEGSL Zou et al. (2023), DisamGCL Zhao et al. (2024) with more details provided in Appendix D.
Implementation Details. We adopt Vicuna-v1.5-7B Zheng et al. (2024) as the LLM model and the Bloom model Le Scao et al. (2023) with 560M and 1B parameters as SLMs. Consistent with most existing works Zhu et al. (2020), we randomly split the nodes into train, validation and test sets with a proportion of 48%/32%/20% for Cornell, Texas, Wisconsin and Actor datasets. We set the proportion of Amazon dataset as 50%/25%/25% following Platonov et al. (2023). In our main experiments, we use FAGCN Bo et al. (2021) as the GNN backbone, adhering to the parameter settings outlined in FAGCN. We only fine-tune the other hyperparameters based on the performance observed on the validation sets. All experiments are repeated 10 times, and we report the averaged results with standard deviation. To ensure a fair comparison, we use the initial node features derived from the Vicuna 7B model for all methods. More implementation details and hyper-parameters are introduced in Appendix B and C.